OpenStudy (anonymous):

For questions 1& 2 write each expression using a single exponent. 1. x^9/ x^2 a) x^11 b) x^7 c) 7 d) x^5 2. h^14/h^2 a) h^-9 b) x^7 c) h^19 d) h^9 Thank you

OpenStudy (anonymous):

\[\frac{x^9}{x^2} = x^9 \cdot \frac{1}{x^2} = x^9 \cdot x^{-2}\]

OpenStudy (anonymous):

now, with \[x^{a} * x^{b} = x^{a+b}\]can you deduce what the result will be?

OpenStudy (anonymous):

Your writing is so tiny c; umm, at the end is x^9 *times* x^-2.. am I right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you sure that you got the possible answers for 2. right? Because the correct one isn't there

OpenStudy (anonymous):

Yeah, those are the only answers there are "/ Maybe it is x^7 ?

OpenStudy (anonymous):

yes, that's correct. And for 2.) it should be h^12, but that isn't there actually

OpenStudy (anonymous):

Oops, I think for question 2 I put answer b) wrong, it is actually 1/h^9

OpenStudy (anonymous):

no, that would be equivalent to h^-9. h^12 should be somewhere there, trust me, h^14 / h^2 = h^12

OpenStudy (anonymous):

Oh alright, well thank you :)